# REST: Diffusion-based Real-time End-to-end Streaming Talking Head Generation via ID-Context Caching and Asynchronous Streaming Distillation

### v1

@@ Abstract
Diffusion models have significantly advanced the field of talking head generation. However, the slow inference speeds and non-autoregressive paradigms severely constrain the application of diffusion-based THG models. In this study, we propose REST, the first diffusion-based, real-time, end-to-end streaming audio-driven talking head generation framework. To support real-time end-to-end generation, a compact video latent space is first learned through high spatiotemporal VAE compression. Additionally, to enable autoregressive streaming within the compact video latent space, we introduce an ID-Context Cache mechanism, which integrates ID-Sink and Context-Cache principles to key-value caching for maintaining temporal consistency and identity coherence during long-time streaming generation. Furthermore, an Asynchronous Streaming Distillation (ASD) training strategy is proposed to mitigate error accumulation in autoregressive generation and enhance temporal consistency, which leverages a non-streaming teacher with an asynchronous noise schedule to supervise the training of the streaming student model. REST bridges the gap between autoregressive and diffusion-based approaches, demonstrating substantial value for applications requiring real-time talking head generation. Experimental results demonstrate that REST outperforms state-of-the-art methods in both generation speed and overall performance.
@@ 1 Introduction
Recent advances in talking head generation have been significantly driven by the introduction of diffusion models (Croitoru et al., 2023). Compared to earlier approaches (Prajwal et al., 2020), diffusion-based THG methods (Wang et al., 2025b, c) demonstrate substantial improvements in both lip-sync accuracy and visual naturalness. Despite the improvements, two major limitations remain in current diffusion-based THG systems. First, end-to-end diffusion models suffer from slow inference speeds, typically requiring tens to hundreds of seconds to generate a five-second video (Meng et al., 2025). Second, most existing diffusion-based models operate in a non-streaming manner, which further increases latency for end users (Wang et al., 2025b). Conversely, autoregressive (AR) THG methods (Chu et al., 2025) typically generate intermediate facial motion representations from speech in a sequential manner, which are then rendered into video via a separate rendering module, achieving lower latency and supporting streaming inference beyond diffusion-based methods. However, the output video quality of two-stage AR methods is often limited by the expressiveness of the generated motion representations, resulting in generally inferior performance compared to end-to-end diffusion-based approaches.
To bridge the gap between diffusion-based and AR-based THG methods in addressing key challenges in the field of THG, in this research we introduce REST, the first diffusion-based real-time end-to-end streaming audio-driven talking head generation framework. We first utilize a Variational Autoencoder (VAE) with high spatiotemporal compression to derive a compact latent video representation, reducing the computational burden of end-to-end diffusion. Inspired by AR methods, we integrate a novel ID-Context Cache mechanism into our diffusion transformer backbone. By combining ID-Sink and Context-Cache, ID-Context Cache enables chunk-by-chunk AR inference while enhancing identity fidelity and inter-frame temporal consistency. Furthermore, to enhance temporal smoothness and generation quality in streaming generation, an Asynchronous Streaming Distillation (ASD) scheme is proposed, which leverages latent spatiotemporal context generated by a non-streaming teacher model to supervise the streaming student model.
@@
  • We propose REST, the first diffusion-based real-time end-to-end streaming talking head generation model that bridges the gap between diffusion-based and AR-based THG methods.
  • We introduce a novel ID-Context Cache mechanism that incorporates ID-Sink and Context KV Cache principles, enabling AR-style inference while maintaining identity fidelity and inter-frame temporal consistency.
  • We further propose an Asynchronous Streaming Distillation scheme for streaming teacher-student supervised learning, leveraging non-streaming spatiotemporal latent context to supervise the streaming student model, achieving better streaming performance.
@@ 2.1 Audio-driven Talking Head Generation
Audio-driven Talking Head Generation (THG) is a cross-modal audio-to-video synthesis task aimed at generating realistic talking head videos conditioned solely on a reference image and speech audio input. Early efforts in Audio-driven THG (Prajwal et al., 2020; Stypułkowski et al., 2024) primarily focused on achieving lip-sync between the generated video and the input speech. Subsequent work (Wang et al., 2021; Zhang et al., 2023) expanded beyond lip-sync to improve the overall naturalness of the output, regrading natural facial expressions and motions. Recently, the success of diffusion models (Nichol & Dhariwal, 2021; Song et al., 2020) in image and video generation (Croitoru et al., 2023) has revolutionized THG. End-to-end THG frameworks built upon pretrained image or video diffusion models have demonstrated superior performance on lip-sync and motion naturalness. However, these improvements come at the cost of substantially increased inference latency. Moreover, to ensure better temporal consistency, most current diffusion-based THG models typically adopt non-autoregressive paradigms that process long video chunks. As a result, users must wait for the entire chunk to be synthesized before viewing, which hinders real-world deployment. To address this limitation, we propose an end-to-end streaming THG framework based on ID-Context Cache, along with a novel streaming teacher-student learning pipeline that achieves identity fidelity and temporal consistency comparable to non-streaming models.
@@ 2.2 Fast Talking Head Generation
Talking Head Generation differs from general video generation in its stronger demand for real-time performance, particularly in applications such as human-computer interaction and game production (Zhen et al., 2023). Conventional efficient THG often adopts a two-stage generation paradigm. In this approach, low-dimensional 2D or 3D motion representations (Li et al., 2025; Liu et al., 2024), are utilized as video driving signals, and a renderer is employed to generate videos from these representations, accelerating inference by converting the high-dimensional latent generation task into a low-dimensional motion generation problem. Motion representation generation is primarily approached via diffusion or AR schemes. Motion-space diffusion methods such as AniTalker (Liu et al., 2024) and Ditto (Li et al., 2025) utilize lightweight diffusion models to predict motion representations, while AR-based approaches (Chu et al., 2025) generate motions in a context-aware manner to enable streaming capabilities. However, two-stage methods are bottlenecked by the fidelity of generated motion representations, generally yielding naturalness inferior to end-to-end approaches (Wang et al., 2025a; Ji et al., 2025). READ (Wang et al., 2025a) represents the first attempt at real-time end-to-end talking head generation, achieving audio-visual alignment in a highly compressed latent space, but its non-autoregressive design still fails to support streaming real-time synthesis. In this work, we propose the first diffusion-based end-to-end streaming THG model that successfully bridges AR and diffusion paradigms.
@@ 3 Methods
The total framework of REST is shown in Fig. 2. Sec. 3.1 outlines the necessary preliminaries relevant to our work. Sec. 3.2 details the proposed model architecture with ID-Context Cache. Sec. 3.3 focuses on the model training methodology guided by ASD. And the final section gives an introduction on the AR-style inference methodology of the parametrized streaming model.
@@ 3.1 Preliminaries
Task Definition. Audio-driven THG task can be formulated as a mapping from a reference image $\mathit{\boldsymbol{I}_{\text{ref}}}\in\mathbb{R}^{H\times W\times 3}$ and speech $\mathit{\boldsymbol{S}_{1:F_{a}}}$ to an output video $\mathit{\boldsymbol{X}}\in\mathbb{R}^{H\times W\times 3\times F_{v}}$, where $F_{a}$ and $F_{v}$ denotes the length of speech and output video.
@@
where ${\mathit{\boldsymbol{u}}}({\mathit{\boldsymbol{Z}}}(t),t)$ is the target gound-truth vector field.
@@ 3.2 Streaming Diffusion with ID-Context Cache
where $H_{w}$ and $h_{w}$ indicates the window sizes, $D_{\text{A}}$ and $d_{\text{A}}$ denotes the hidden dims.
ID-Context Cache for autoregressive diffusion. To enable autoregressive, context-aware generation in audio-to-video diffusion, we introduce a novel DiT-based backbone that implements the proposed ID-Context Cache mechanism. Our Streaming A2V-DiT framework comprises 28 transformer blocks, each integrating three core modules: Self-Attention with ID-Context Cache, 3D Full-Attention for text conditioning, and Frame-level 2D Cross-Attention for audio conditioning. Theoretically, the ID-Context Cache reconfigures the attention topology to enable chunk-by-chunk autoregressive inference, simultaneously enforcing ID fidelity and temporal causality. This is realized through the two following theoretically motivated components.
ID-Sink Principle: ID-Sink models the Key-Value (KV) embeddings of the reference image as persistent sinks across all chunks, establishing a global semantic anchor to preserve identity fidelity. Context-Cache Principle: Context Cache approximates a continuous temporal flow by extending the effective temporal receptive field. It reconstructs local temporal dependencies by concatenating the KV embeddings of the preceding chunk with the current KV embedding during self-attention, thereby mitigating boundary discontinuities.
@@
where $i$ denotes the $i$-th chunk and $j$ denotes the $j$-th block. Then, the causal dependency chain is reconstructed by concatenating the static ID anchor with the dynamically updated contextual cache from the preceding chunk, following the proposed ID-Sink and Context Cache principles:
@@
The self-attention is then operated based on the refined causal dependency chain, as follows:
@@
Regarding cross-modal audio conditioning, the speech latent code ${\boldsymbol{E}}=[\mathit{\boldsymbol{e}}_{0}\mathbin{\|}\mathit{\boldsymbol{e}}_{1}\mathbin{\|}\cdots\mathbin{\|}\mathit{\boldsymbol{e}}_{k}]$ temporally aligned with the video chunks. By enforcing a frame-level spatial cross-attention with a temporal receptive field of chunk length, we achieve precise streaming audio-visual alignment within the compact latent space, formulated as:
@@
where $\mathit{\boldsymbol{H}}_{ij},\mathit{\boldsymbol{H}}_{ij}^{\text{A}}\in\mathbb{R}^{h\times w\times f\times d}$ denotes the unpatchified hidden states before and after frame-level spatial cross-attention of $i\!-\!\text{th}$ attention block. Our proposed design enables the AR-style streaming generation of video latent chunks that are synchronized with the corresponding speech conditions.
@@ 3.3 Training with Asynchronous Streaming Distillation
To mitigate autoregressive error propagation and enforce inter-chunk temporal consistency in chunk-by-chunk streaming diffusion, we introduce the Asynchronous Streaming Distillation (ASD) paradigm. Theoretically, ASD aims to distill global temporal priors from a non-streaming teacher into a streaming student through asynchronous noise conditioning.
@@
$\!{\mathit{\boldsymbol{Z}}}(\boldsymbol{t})\!=\!(\boldsymbol{1}-\boldsymbol{t})\odot{\mathit{\boldsymbol{Z}}}(0)+\boldsymbol{t}\odot\boldsymbol{\epsilon}$ (16)
@@
$\boldsymbol{v}$ $=\boldsymbol{\epsilon}-Z(0)$ (17)
@@
where $\boldsymbol{v}$ denotes the correct vector field under the Gaussian probability path of the asynchronous add-noise process.
Streaming Student Training. Subsequently, to emulate inference-time causal dynamics, the student model is trained via a sequential chunk-wise paradigm. We apply the same add-noise paradigm as teacher model to obtain noisy latents ${\mathit{\boldsymbol{Z}}}(t)$ from the initial latents ${\mathit{\boldsymbol{Z}}}(0)$, which are then partitioned into chunks $\widetilde{\mathit{\boldsymbol{Z}}}(t)=[\mathit{\boldsymbol{z}}_{\text{R}}\mathbin{\|}\mathit{\boldsymbol{z}}_{1}(t)\mathbin{\|}\cdots\mathbin{\|}\mathit{\boldsymbol{z}}_{k}(t)]$. These chunks are processed sequentially by the student model with the proposed ID-Context Cache mechanism to execute autoregressive generation. Also, the final training objective of the student network parameters $\mathcal{S_{\theta}}$ is as follows, conditioned on the audio latent chunks $\widetilde{\boldsymbol{E}}=[\mathit{\boldsymbol{e}}_{0}\mathbin{\|}\mathit{\boldsymbol{e}}_{1}\mathbin{\|}\cdots\mathbin{\|}\mathit{\boldsymbol{e}}_{k}]$:
$\boldsymbol{v}$ $=\boldsymbol{\epsilon}-Z(0)$ (19)
@@
Teacher-Student Learning with ASD. The non-streaming teacher model leverages global attention to achieve superior temporal consistency, while the streaming student is limited by local causal attention under streaming constraints. Simply enlarging the Context Cache to broaden the receptive field leads to unsustainable computational costs. To address this fundamental limitation, we propose distribution matching between the non-streaming teacher and the streaming student, introducing global consistency constraints while adhering to streaming constraints. Specifically, we introduce a contrastive distribution loss that minimizes the distribution distance between teacher output flow $\hat{\mathit{\boldsymbol{v}}}=[\hat{\mathit{{v}}}_{0},\cdots,\hat{\mathit{{v}}}_{f}]$ and student output flow $\hat{\widetilde{\mathit{\boldsymbol{v}}}}=[\hat{\widetilde{\mathit{{v}}}}_{0},\cdots,\hat{\widetilde{\mathit{{v}}}}_{f}]$ at corresponding spatiotemporal positions, while maximizing the distance at different positions, as follows:
$\mathcal{L}_{\text{CON}}\!=\!-\frac{1}{f}\sum_{i=1}^{f}\mathrm{log}\!\left(\frac{\mathrm{exp}\left(\frac{\mathrm{sim}(\hat{\widetilde{\mathit{\boldsymbol{v}}}}_{i},\hat{\mathit{\boldsymbol{v}}}_{i})}{\tau}\right)}{\sum_{j=1,j\neq i}^{f}\mathrm{exp}\left(\frac{\mathrm{sim}(\hat{\widetilde{\mathit{\boldsymbol{v}}}}_{i},{\hat{\mathit{\boldsymbol{v}}}}_{j})}{\tau}\right)}\right)$ (21)
Furthermore, to enforce inter-chunk temporal consistency, we introduce a smoothness matching objective that minimizes the divergence in second-order optical flow between the outputs of teacher and student models at corresponding temporal positions. This regularization serves as a global consistency constraint, formulated as:
$\Delta v_{i}$ $=v_{i+1}-2v_{i}+v_{i-1}$ (22)
$\mathcal{L}_{\text{SMO}}$ $=\sum\nolimits_{i=1}^{f-1}\Delta v_{i}$ (23)
@@
$\mathcal{L}_{\text{SD}}=\mathcal{L}_{\mathcal{S}_{\theta}}+\mathcal{L}_{\text{CON}}+\mathcal{L}_{\text{SMO}}$ (24)
By aligning asynchronous teacher-student learning with the proposed constraints, global contextual knowledge from the non-streaming teacher can be effectively distilled into the streaming student, thereby enhancing the stability and temporal consistency in streaming generation.

### v2

@@ Abstract
Diffusion models have significantly advanced the field of talking head generation (THG). However, slow inference speeds and prevalent non-autoregressive paradigms severely constrain the application of diffusion-based THG models. In this study, we propose REST, a pioneering diffusion-based, real-time, end-to-end streaming audio-driven talking head generation framework. To support real-time end-to-end generation, a compact video latent space is first learned through a spatiotemporal variational autoencoder with a high compression ratio. Additionally, to enable semi-autoregressive streaming within the compact video latent space, we introduce an ID-Context Cache mechanism, which integrates ID-Sink and Context-Cache principles into key-value caching for maintaining identity consistency and temporal coherence during long-term streaming generation. Furthermore, an Asynchronous Streaming Distillation (ASD) strategy is proposed to mitigate error accumulation and enhance temporal consistency in streaming generation, leveraging a non-streaming teacher with an asynchronous noise schedule to supervise the streaming student. REST bridges the gap between autoregressive and diffusion-based approaches, achieving a breakthrough in efficiency for applications requiring real-time THG. Experimental results demonstrate that REST outperforms state-of-the-art methods in both generation speed and overall performance.
@@ 1 Introduction
Recent advances in talking head generation have been significantly driven by the introduction of diffusion models (Croitoru et al., 2023). Diffusion-based THG methods (Wang et al., 2025b, c) demonstrate substantial improvements in both lip-sync accuracy and visual naturalness. Despite the improvements, two major limitations remain in current diffusion-based THG systems. First, end-to-end diffusion models suffer from slow inference speeds, typically requiring tens to hundreds of seconds to generate a five-second video (Meng et al., 2025). Second, most existing diffusion-based models operate in a non-streaming manner, which further increases latency for end users (Wang et al., 2025b). Conversely, autoregressive (AR) THG methods (Chu et al., 2025) typically generate intermediate facial motion representations from speech in a sequential manner, which are then rendered into video via a separate rendering module, achieving lower latency and supporting streaming inference beyond diffusion-based methods. However, the output video quality of two-stage AR methods is limited by the expressiveness of the generated motion representations, resulting in generally inferior performance compared to end-to-end diffusion-based approaches (Wang et al., 2025a).
To bridge the gap between diffusion-based and AR-based THG methods in addressing key challenges in the field of THG, in this research we introduce a novel diffusion-based Real-time End-to-end Streaming Talking head generation (REST) framework. We first utilize a variational autoencoder (VAE) with high spatiotemporal compression to learn a compact latent space, reducing the computational burden of end-to-end diffusion. Inspired by AR methods, we integrate a novel ID-Context Cache mechanism into our diffusion transformer backbone. By combining ID-Sink and Context-Cache, ID-Context Cache enables streaming inference while enhancing identity and temporal consistency. Furthermore, to address the quality degradation challenge inherent in streaming generation, an Asynchronous Streaming Distillation (ASD) scheme is proposed, which leverages latent spatiotemporal context generated by a non-streaming teacher model to supervise the streaming student model.
@@
  • We propose REST, a novel framework for real-time, end-to-end, streaming talking head generation that effectively bridges the gap between diffusion and AR paradigms. To the best of our knowledge, REST is the first end-to-end method to achieve real-time, streaming audio-driven talking head diffusion on a single GPU.
  • We introduce a novel ID-Context Cache mechanism to facilitate high-quality streaming generation. By integrating ID-Sink and Context-Cache principles, our approach maintains high ID fidelity and temporal coherence across frames under streaming constraints.
  • We further propose an Asynchronous Streaming Distillation (ASD) scheme to bridge the performance gap between streaming and non-streaming models by integrating information-theoretic alignment and kinematic constraints into a unified distillation objective.
@@ 2.1 Audio-driven Talking Head Generation
Audio-driven Talking Head Generation (THG) is a cross-modal synthesis task aimed at generating a realistic talking head video conditioned on a reference image and a speech signal. Early efforts (Prajwal et al., 2020; Stypułkowski et al., 2024) primarily focused on achieving lip-sync between the generated video and the input speech. Subsequent work (Wang et al., 2021; Zhang et al., 2023) expanded beyond lip-sync to improve the overall naturalness of the output, regarding natural facial expressions and motions. Recently, the success of diffusion models (Nichol and Dhariwal, 2021; Song et al., 2020) in image and video generation (Croitoru et al., 2023) has revolutionized THG. End-to-end THG frameworks built upon pretrained diffusion models have demonstrated superior performance on lip-sync and naturalness. However, these improvements come at the cost of substantially increased latency (Wang et al., 2025a). Moreover, to ensure better temporal consistency, most current diffusion-based THG models typically adopt non-autoregressive paradigms (Ji et al., 2025). As a result, users must wait for the entire video to be synthesized before viewing, which hinders real-world deployment. To address this limitation, we propose an end-to-end streaming THG framework based on ID-Context Cache and a streaming teacher-student learning scheme, achieving robust streaming generation with superior ID and temporal consistency.
@@ 2.2 Fast Talking Head Generation
Talking Head Generation differs from general video generation in its stronger demand for real-time performance, particularly in applications such as human-computer interaction and game production (Zhen et al., 2023). Conventional efficient THG often adopts a two-stage generation paradigm. In this approach, low-dimensional 2D or 3D motion representations (Li et al., 2025; Liu et al., 2024), are utilized as video driving signals, and a renderer is employed to generate videos from these representations, accelerating inference by converting the high-dimensional latent generation task into a low-dimensional motion generation problem. Motion representation generation is primarily approached via diffusion or AR schemes. Motion-space diffusion methods such as AniTalker (Liu et al., 2024) and Ditto (Li et al., 2025) utilize lightweight diffusion models to predict motion representations, while AR-based approaches (Chu et al., 2025; Sung-Bin et al., 2024) generate motions in a context-aware manner to enable streaming capabilities. However, two-stage methods are bottlenecked by the fidelity of generated motion representations, generally yielding naturalness inferior to end-to-end approaches (Ji et al., 2025; Meng et al., 2025). READ (Wang et al., 2025a) represents the first attempt at real-time end-to-end talking head generation, achieving audio-visual alignment in a compressed latent space, but its non-autoregressive design still fails to support streaming real-time synthesis. In this work, we propose a pioneering diffusion-based end-to-end streaming THG model that successfully bridges AR and diffusion paradigms.
@@ 3 Methods
The total framework of REST is shown in Fig. 2. Sec. 3.1 outlines the necessary preliminaries relevant to our work. Sec. 3.2 details the proposed model architecture of REST with ID-Context Cache. Sec. 3.3 focuses on the model training methodology guided by the proposed ASD. And the final section gives an introduction to the AR-style inference methodology of the parametrized streaming REST model.
@@ 3.1 Preliminaries
Task Definition. Audio-driven THG task can be formulated as a mapping from a reference image $\mathit{\boldsymbol{I}_{\text{ref}}}\in\mathbb{R}^{H\times W\times 3}$ and speech $\mathit{\boldsymbol{S}_{1:F_{a}}}$ to an output video $\mathit{\boldsymbol{X}}\in\mathbb{R}^{H\times W\times F\times D_{\text{v}}}$ ($D_{\text{v}}=3$), $F_{a}$ and $F$ denote the length of speech and video.
@@
where ${\mathit{\boldsymbol{u}}}({\mathit{\boldsymbol{Z}}}(t),t)$ is the target ground-truth vector field.
@@ 3.2 Streaming Diffusion with ID-Context Cache
where $H_{w}$ and $h_{w}$ indicates the window sizes, $D_{\text{A}}$ and $d_{\text{A}}$ denotes the hidden dims. Detailed architecture information is provided in Sec. B.2 of the Appendix.
ID-Context Cache for semi-autoregressive diffusion. To enable semi-autoregressive, context-aware generation in audio-to-video diffusion, we introduce a novel DiT-based backbone that implements the proposed ID-Context Cache mechanism. Our Streaming A2V-DiT framework comprises 28 transformer blocks, each integrating three core modules: Self-Attention with ID-Context Cache, 3D Full-Attention for text conditioning, and Frame-level 2D Cross-Attention for audio conditioning (please refer to Appendix Sec. B.2 for details). Theoretically, the ID-Context Cache reconfigures the attention topology to enable chunk-by-chunk semi-autoregressive inference, simultaneously enforcing ID fidelity and temporal consistency. This is realized through the two following theoretically motivated components.
ID-Sink Principle: ID-Sink models the Key-Value (KV) embeddings of the reference image as a persistent sink across all chunks, establishing a global semantic anchor to preserve identity fidelity. Context-Cache Principle: Context Cache approximates a continuous temporal flow by extending the effective temporal receptive field. It reconstructs local temporal dependencies by concatenating the KV embeddings of the preceding chunk with the current KV embedding during self-attention, thereby mitigating boundary discontinuities of streaming generation.
@@
where $i$ denotes the $i$-th chunk and $j$ denotes the $j$-th block of A2V-DiT. Then, the causal chain is reconstructed by concatenating the static ID anchor with the dynamically updated contextual cache from the preceding chunk, following the proposed ID-Sink and Context-Cache principles:
@@
The self-attention is then operated on hidden states $\mathit{\boldsymbol{H}}_{ij}$ based on the refined causal dependency chain, as follows:
@@
Regarding cross-modal audio conditioning, the speech latent code $\boldsymbol{E}=[\boldsymbol{e}_{0}\mathbin{\|}\boldsymbol{e}_{1}\mathbin{\|}\cdots\mathbin{\|}\boldsymbol{e}_{k}]$ ($k$ denotes the number of chunks), which is temporally aligned with the video chunks, serves as the driving condition. By enforcing a frame-level spatial cross-attention with a temporal receptive field of chunk length, we achieve precise streaming audio-visual alignment within the compact latent space, formulated as:
@@
where $\mathit{\boldsymbol{H}}_{ij},\mathit{\boldsymbol{H}}_{ij}^{\text{A}}\in\mathbb{R}^{h\times w\times f\times d}$ denotes the hidden states before and after frame-level spatial cross-attention of $i\!-\!\text{th}$ attention block. Our proposed design enables the semi-AR streaming generation of video latent chunks that are synchronized with the corresponding speech conditions. Additional analysis of the ID-Context Cache is provided in Sec. B.1.
@@ 3.3 Training with Asynchronous Streaming Distillation
To mitigate autoregressive error propagation and enforce generation quality in chunk-by-chunk streaming diffusion, we introduce the Asynchronous Streaming Distillation (ASD) scheme. Theoretically, ASD aims to distill global temporal priors from a non-streaming teacher into a streaming student through asynchronous noise conditioning.
@@
$\!{\mathit{\boldsymbol{Z}^{\mathcal{T}}}}(\boldsymbol{t})\!=\!(\boldsymbol{1}-\boldsymbol{t})\odot{\mathit{\boldsymbol{Z}}}(0)+\boldsymbol{t}\odot\boldsymbol{\epsilon}$ (16)
@@
$\boldsymbol{v}^{\mathcal{T}}$ $=\boldsymbol{\epsilon}-{\mathit{\boldsymbol{Z}^{\mathcal{T}}}}(0)$ (17)
@@
where $\boldsymbol{v}^{\mathcal{T}}$ denotes the correct velocity flow under the Gaussian probability path of the asynchronous add-noise process.
Streaming Student Training. Subsequently, to emulate inference-time causal dynamics, the student model is trained via a sequential chunk-wise paradigm. We apply the same add-noise paradigm as teacher model to obtain noisy latents $\boldsymbol{Z}^{\mathcal{S}}(\boldsymbol{t})$ from the initial latents $\boldsymbol{Z}(0)$, which are then partitioned into chunks $\boldsymbol{Z}^{\mathcal{S}}(\boldsymbol{t})=[\mathit{\boldsymbol{z}}_{\text{R}}\mathbin{\|}\mathit{\boldsymbol{Z}}_{1}(t)\mathbin{\|}\cdots\mathbin{\|}\mathit{\boldsymbol{Z}}_{k}(t)]$. These chunks are processed sequentially by the student model with the proposed ID-Context Cache mechanism to execute semi-autoregressive generation. Also, the final training objective of the student network parameters $\mathcal{S_{\theta}}$ is as follows, conditioned on the audio latent chunks $\boldsymbol{E}=[\boldsymbol{E}_{1}\mathbin{\|}\cdots\mathbin{\|}\boldsymbol{E}_{k}]$:
$\boldsymbol{v}^{\mathcal{S}}$ $=\boldsymbol{\epsilon}-\boldsymbol{Z}(0)$ (19)
@@
Teacher-Student Learning with ASD. There exists a fundamental information asymmetry between the teacher and student. The non-streaming teacher model leverages global attention to achieve superior consistency, while the streaming student is limited by local causal attention under streaming constraints. To address this fundamental limitation, we propose ASD between the non-streaming teacher and the streaming student. From an information-theoretic perspective, we aim to maximize the average mutual information (MI) between the generative distributions of student and teacher, thereby introducing global consistency constraints under streaming conditions. This is achieved by maximizing the MI variational lower bound via frame-wise contrastive learning, specifically by minimizing the InfoNCE loss (Oord et al., 2018) of the output velocity flows ${\mathit{\boldsymbol{v}^{\mathcal{T}}}}=\{{\mathit{\boldsymbol{v}}}^{\mathcal{T}}_{i}\}_{i=0}^{f}$ and ${\mathit{\boldsymbol{v}^{\mathcal{S}}}}=\{{\mathit{\boldsymbol{v}}}^{\mathcal{S}}_{i}\}_{i=0}^{f}$ (proof in Appendix Sec. C.1). The contrastive objective can be formulated as follows:
$\mathcal{L}_{\text{CON}}=-\frac{1}{f}\sum_{i=1}^{f}\log\left(\frac{\exp\left(\mathrm{sim}(\boldsymbol{v}^{\mathcal{S}}_{i},\boldsymbol{v}^{\mathcal{T}}_{i})/\tau\right)}{\sum\limits_{j=1}^{f}\exp\left(\mathrm{sim}(\boldsymbol{v}^{\mathcal{S}}_{i},\boldsymbol{v}^{\mathcal{T}}_{j})/\tau\right)}\right)$ (21)
Furthermore, from a kinematic perspective, to enforce inter-chunk temporal consistency, we introduce a smoothness matching objective that minimizes the divergence in second-order optical flow between the outputs of teacher and student to serve as a global consistency constraint, formulated as:
$\Delta^{2}\boldsymbol{v}_{i}$ $=\boldsymbol{v}_{i+1}-2\boldsymbol{v}_{i}+\boldsymbol{v}_{i-1}$ (22)
$\mathcal{L}_{\text{SMO}}$ $=\frac{1}{f-1}\sum\nolimits_{i=1}^{f-1}(\Delta^{2}\boldsymbol{v}^{\mathcal{S}}_{i}-\Delta^{2}\boldsymbol{v}^{\mathcal{T}}_{i})^{2}$ (23)
@@
$\mathcal{L}_{\text{ASD}}=\mathcal{L}_{\mathcal{S}_{\theta}}+\alpha\mathcal{L}_{\text{CON}}+\beta\mathcal{L}_{\text{SMO}}$ (24)
Analysis of the loss complementarity is detailed in Sec. C.3. ASD serves to transfer global contextual knowledge from a non-streaming teacher to the streaming student, improving the stability and consistency of streaming generation.